# LongReason: A Synthetic Long-Context Reasoning Benchmark via Context Expansion

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable progress in understanding long-context inputs. However, benchmarks for evaluating the long-context reasoning abilities of LLMs fall behind the pace. Existing benchmarks often focus on a narrow range of tasks or those that do not demand complex reasoning. To address this gap and enable a more comprehensive evaluation of the long-context reasoning capabilities of current LLMs, we propose a new synthetic benchmark, LongReason, which is constructed by synthesizing long-context reasoning questions from a varied set of short-context reasoning questions through context expansion. LongReason consists of 794 multiple-choice reasoning questions with diverse reasoning patterns across three task categories: reading comprehension, logical inference, and mathematical word problems. We evaluate 21 LLMs on LongReason, revealing that most models experience significant performance drops as context length increases. Our further analysis shows that even state-of-the-art LLMs still have significant room for improvement in providing robust reasoning across different tasks. We have open-sourced LongReason [^1] https://huggingface.co/datasets/lz1bytedance/LongReason to support the comprehensive evaluation of LLMs’ long-context reasoning capabilities.
LongReason: A Synthetic Long-Context Reasoning Benchmark via Context Expansion
Zhan Ling 1  Kang Liu 2  Kai Yan 3  Yifan Yang 2  Weijian Lin 2 Ting-Han Fan 2  Lingfeng Shen 2  Zhengyin Du 2  Jiecao Chen 2 1 UC San Diego  2 Bytedance  3 UIUC
@@ 1 Introduction
ZeroSCROLLS Shaham et al. (2023) $\sim$10K
L-Eval An et al. (2024) $\sim$8K
BAMBOO Dong et al. (2023a) $\sim$16K
LongBench Bai et al. (2023) $\sim$8K
LooGLE Li et al. (2023) $\sim$20K
InfiniteBench Zhang et al. (2024a) $\sim$200K
Loong Wang et al. (2024) $\sim$250K
Needle-in-a-haystack Kamradt (2023) any
RULER Hsieh et al. (2024) any
@@
In recent years, large language models (LLMs) OpenAI: Josh Achiam et al. (2023); Reid et al. (2024b); Anthropic (2024b); Meta.AI (2024); Jiang et al. (2023); Team (2024) have demonstrated remarkable advances in diverse natural language processing tasks. The ability to comprehend and reason over long inputs is essential for downstream applications, including multi-turn conversations Tan et al. (2024), document understanding Masry and Hajian (2024) retrieval-augmented generation Yao et al. (2022); Xu et al. (2023), and language agents Zhao et al. (2024); Zhang et al. (2024b). Meanwhile, extensive efforts in deep learning system Dao et al. (2022); Dao (2023); Chen et al. (2023b); Ratner et al. (2022) research have been devoted to optimizing computational overhead to support increasing numbers of input tokens, which has led to growing attention on long-context LLMs. Now, both proprietary and open-source LLMs can support up to millions of input tokens Reid et al. (2024b); Mistral.AI (2024b); GLM et al. (2024).
However, despite the rapid development of long-context language models, benchmarks have lagged behind. One of the key challenges is dataset construction, as long-context question-answering data is relatively scarce on the internet. To address this, prevalent long-context benchmarks have utilized synthetic tasks like passkey retrieval Mohtashami and Jaggi (2023), needle-in-a-haystack (NIAH) Kamradt (2023); Zhao et al. (2024), and variable tracking Hsieh et al. (2024) to evaluate long-context LLMs. However, these tasks are often unrealistic and involve reasoning processes that differ significantly from those in real-world applications. Alternatively, some research efforts have involved human annotation of realistic questions and gold answers over one or multiple long documents Dong et al. (2023b); Wang et al. (2024); Li et al. (2023). However, creating realistic long-context tasks from extensive texts is both challenging and time-consuming, even for human experts Wang et al. (2024). This limitation restricts the expansion of datasets to accommodate arbitrary context lengths and the ability to support controllable context. As shown in Table 1, existing benchmarks either rely on a limited number of synthetic tasks, demand significant human effort to read long contexts, or lacking controllable contexts and support for arbitrary context lengths. Furthermore, existing datasets Dong et al. (2023b); Wang et al. (2024); Li et al. (2023) often utilize documents from specific domains, such as financial reports or legal cases, as input, which can inherently limit the diversity of task categories. Consequently, they tend to focus on a narrow set of tasks, such as comparison or classification, rather than evaluating more complex and challenging tasks that require chain-of-thought reasoning.
@@ 2 Related Work
Long-Context Large Language Models Recent advancements in deep learning system have significantly propelled the development of long-context large language models (LLMs). One of the key challenges in scaling these models is the quadratic time and space complexity inherent in computing self-attention over long sequences. To mitigate this computational burden, efficient self-attention algorithms Dao et al. (2022); Dao (2023); Liu et al. (2023) have been introduced, reducing memory overhead, and novel training methods Li et al. (2021); Chen et al. (2023b) facilitate the training of these long-context models. As Rotary Position Embedding (RoPE) Su et al. (2024) is widely used for positional encoding in many open-source models Meta.AI (2024); Team (2024); Mistral.AI (2024a), recent research Chen et al. (2023a); Xiong et al. (2023); Peng et al. (2023b); Liu et al. (2024); Ding et al. (2024); Zhu et al. (2024) has focused on adapting RoPE from pre-trained short-context models to effectively handle longer sequences. Moreover, new architectures Gu and Dao (2023); Peng et al. (2023a); Bulatov et al. (2022, 2023); Botev et al. (2024) have been developed to efficiently process long-context inputs. Consequently, state-of-the-art language models OpenAI (2024b, a); Reid et al. (2024a); Anthropic (2024a); Meta.AI (2024); Mistral.AI (2024a); Team (2024); GLM et al. (2024) now support context windows ranging from 128K to millions of tokens, enabling the exploration of reasoning abilities over extensive contexts with LLMs.
@@
Long-Context Benchmarks As the context window of current LLMs expands rapidly, numerous benchmarks have been proposed to evaluate their capabilities. In early benchmarks such as ZeroSCROLLS Shaham et al. (2023), L-Eval An et al. (2024), BAMBOO Dong et al. (2023b), LongBench Bai et al. (2023), and LooGLE Li et al. (2023), the average input length remains under 25K tokens, which is far shorter than the context window size supported by existing LLMs. Recently, some research has begun to explore using synthetic datasets, which can support controllable context lengths, to evaluate the long-context abilities of LLMs. Needle-in-a-Haystack and its variants Kamradt (2023); Zhao et al. (2024) primarily evaluate retrieval abilities by inserting relevant information into extensive irrelevant corpora and testing the LLMs’ capacity to extract it. Additionally, RULER Hsieh et al. (2024) constructs synthetic tasks based on code-like flexible configurations to assess LLM performance over long contexts. While synthetic tasks can support the evaluation of arbitrarily long contexts, they are limited in scope, focusing on a narrow set of tasks and failing to comprehensively evaluate the reasoning abilities of LLMs in realistic scenarios. Other benchmarks like InfiniteBench Zhang et al. (2024a) and Loong Wang et al. (2024) use human annotations to create questions from given long texts, which contain more diverse tasks but are both time-consuming and costly. Our proposed benchmark, LongReason, focuses on evaluating the long-context reasoning abilities of LLMs, which are created automatically from short reasoning questions without heavy human effort in reading the long context. We conduct a detailed comparison with existing benchmarks in Table 1.
@@ 3.1 Long-context Reasoning Question Construction via Context Expansion
Long-Context Reasoning Question Construction Through Context Expansion Finally, we construct the long-context version of each question by embedding each passage $C_{e}^{i}$ from the expanded context $\bar{C}{\text{expanded}}$ at random positions within a set of irrelevant passages $\bar{C}{\text{irrelevant}}$, forming the final long-context reasoning questions. To create $\bar{C}{\text{irrelevant}}$ , we first collect passages from the Pile Gao et al. (2020) and use an LLM to rewrite each passage to minimize stylistic differences between the synthesized background passages and the irrelevant passages. These rewritten passages are then compiled to form the set of irrelevant passages $\bar{C}_{\text{irrelevant}}$. In LongReason, GPT-4 is used for all data synthesis and self-verification. For each question, we evaluate multiple versions of the synthesized question for comparison, including the original question $Q_{\text{short}}$, the expanded version $Q_{\text{expanded}}$, and long-context versions with context lengths ranging from $8K$ to $128K$. Furthermore, similar to NIAH Kamradt (2023), our pipeline is capable of generating reasoning questions with even longer contexts by incorporating additional irrelevant information.
@@ 3.2 The Statistics of LongReason
11footnotetext: Mistral Large 2 generates an empty response when the question context length reaches 128K.
@@ 4.1 Experimental setup
Models & Inference Setup We select a set of representative LLMs that support long context windows, including 6 closed-source models from 3 model families (GPT, Gemini and Claude) and 15 open-source models spanning a wide range of model sizes (3B to 123B) and claimed context lengths (8K to 2M). Detailed information about these models can be found in Appendix A. For open-source models, we utilize vLLM Kwon et al. (2023), which enables efficient KV cache memory management during inference time. All inferences are performed using bfloat16 precision on 8 NVIDIA A100 GPUs with greedy decoding (temperature=0).
Evaluation setup We evaluate all models on LongReason, which comprises 794 questions, each featuring multiple variations, including the original version, expanded versions, and long-context versions with context lengths of 8K, 16K, 32K, 64K, and 128K. Each input is constructed using a predefined zero-shot chain-of-thought template that combines the background context, followed by the corresponding final inquiry. Detailed information about the prompt template is provided in Appendix D. To assess the reasoning performance of the LLMs, we extract the predicted choice by identifying the first character following the phrase “the answer is” and compare it to the ground-truth option for accuracy.
@@ Appendix A Model Information
Model Aligned Size Context Length Huggingface (Wolf et al., 2019) / API
GPT-4o (OpenAI, 2024b) - 128K gpt-4o-2024-08-06
GPT-4o-mini (OpenAI, 2024a) - 128K gpt-4o-mini-2024-07-18
Gemini-1.5-Pro (Reid et al., 2024b) - 2M gemini-1.5-pro-002
Gemini-1.5-Flash (Reid et al., 2024b) - 2M gemini-1.5-flash-002
Claude-3.5-Sonnet(Anthropic, 2024a) - 200K claude-3-5-sonnet-20240620
Claude-3.5-Haiku(Anthropic, 2024a) - 200K claude-3-5-haiku-20241022
Mistral-Large2 (Mistral.AI, 2024a) 123B 128K mistralai/Mistral-Large-Instruct-2407
Mixtral-8$\times$22B (Jiang et al., 2024) 39B/8$\times$22B 64K mistralai/Mixtral-8x22B-Instruct-v0.1
Mistral-Small (Mistral.AI, 2024c) 22B 32K mistralai/Mistral-Small-Instruct-2409
Mistral-Nemo (Mistral.AI, 2024b) 12B 1M mistralai/Mistral-Nemo-Instruct-2407
Mistral-7B (Jiang et al., 2023) 7B 32K mistralai/Mistral-7B-Instruct-v0.3
Llama3.1 (Meta.AI, 2024) 70B 128K meta-llama/Meta-Llama-3.1-70B-Instruct
Llama3.1 (Meta.AI, 2024) 8B 128K meta-llama/Meta-Llama-3.1-8B-Instruct
Qwen2.5 (Team, 2024) 72B 128K Qwen/Qwen2.5-72B-Instruct
Qwen2.5 (Team, 2024) 32B 128K Qwen/Qwen2.5-32B-Instruct
Qwen2.5 (Team, 2024) 14B 128K Qwen/Qwen2.5-14B-Instruct
Qwen2.5 (Team, 2024) 7B 128K Qwen/Qwen2.5-7B-Instruct
Qwen2.5 (Team, 2024) 3B 32K Qwen/Qwen2.5-3B-Instruct
GLM4-9B (GLM et al., 2024) 9B 128K THUDM/glm-4-9b-chat
Phi3.5-MoE (Abdin et al., 2024) 6.6B/16$\times$3.8B 128K microsoft/Phi-3.5-MoE-instruct
Phi3.5-mini (Abdin et al., 2024) 14B 128K microsoft/Phi-3.5-mini-instruct
@@
Appendix B Hyperparameters for LongReason Construction
@@ Appendix D Prompts
Please expand the above material into multiple independent paragraphs with around 200 words in English, while meeting the following
requirements. 1. Ensure that every key piece of information from the material appears in one paragraph of the expanded text. Try to
place the key information in the middle of the expanded paragraphs.
2. The expanded material need to avoid introducing additional knowledge, reasoning, or any content that might influence the answer to
the "Question about the material."
3. The expanded text should clearly relate to the "Question about the material." Please include hints or references to the question within
each paragraph to maintain this connection. However, do not use words like "question" or "query" explicitly in the expanded text.
4. Each paragraph should be a standalone piece of text, comprehensible without needing to refer to other paragraphs. Minimize the
use of pronouns, particularly those referring to other paragraphs.
@@
Please analyze how to appropriately add background information to expand the material into multiple independent paragraphs based on
the given requirements. Additionally, assess whether the provided material can be easily divided into multiple paragraphs. If not please
only provide only one paragraph in the expanded material.
@@
Present the expanded material as a series of independent paragraphs that meet the above requirements. Add the index like "1."
(do not use any format here) at the beginning of each paragraph (starting from 1), use English, and do not include any extra information.
@@ 2. Does the expanded material contain information that will affect the answer to the question?
Please analyze the expanded material and compare it with the original material. Then, combine the question and analyze the three
questions above.
### Does the expanded material contain all the key information from the original material based on the analysis?
@@
### Does the expanded material contain information that will affect the answer to the question based on the analysis?
@@
### Are all the paragraphs in the expanded material are related to the main topic/character of the question based on the analysis?

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable progress in understanding long-context inputs. However, benchmarks for evaluating the long-context reasoning abilities of LLMs fall behind the pace. Existing benchmarks often focus on a narrow range of tasks or those that do not demand complex reasoning. To address this gap and enable a more comprehensive evaluation of the long-context reasoning capabilities of current LLMs, we propose a new synthetic benchmark, LongReason, which is constructed by synthesizing long-context reasoning questions from a varied set of short-context reasoning questions through context expansion. LongReason consists of 794 multiple-choice reasoning questions with diverse reasoning patterns across three task categories: reading comprehension, logical inference, and mathematical word problems. We evaluate 21 LLMs on LongReason, revealing that most models experience significant performance drops as context length increases. Our further analysis shows that even state-of-the-art LLMs still have significant room for improvement in providing robust reasoning across different tasks. We have open-sourced LongReason to support the comprehensive evaluation of LLMs’ long-context reasoning capabilities.
[…]
@@ 1 Introduction
ZeroSCROLLS [zeroscrolls] $\sim$10K
L-Eval [leval] $\sim$8K
BAMBOO [dong2023bamboo] $\sim$16K
LongBench [bai2023longbench] $\sim$8K
LooGLE [li2023loogle] $\sim$20K
InfiniteBench [zhang2024infty] $\sim$200K
Loong [wang2024leave] $\sim$250K
Needle-in-a-haystack [needleinhaystack] any
RULER [hsieh2024ruler] any
@@
In recent years, large language models (LLMs) [openai2023gpt4, gemini, claude3, llama3-1, mistral, qwen2.5] have demonstrated remarkable advances in diverse natural language processing tasks. The ability to comprehend and reason over long inputs is essential for downstream applications, including multi-turn conversations [tan2024peer], document understanding [masry2024longfin] retrieval-augmented generation [yao2022react, xu2023retrieval], and language agents [zhao2024longagent, zhang2024chain]. Meanwhile, extensive efforts in deep learning system [dao2022flashattention, dao2023flashattention, chen2023longlora, ratner2022parallel] research have been devoted to optimizing computational overhead to support increasing numbers of input tokens, which has led to growing attention on long-context LLMs. Now, both proprietary and open-source LLMs can support up to millions of input tokens [gemini, mistralnemo, glm4].
However, despite the rapid development of long-context language models, benchmarks have lagged behind. One of the key challenges is dataset construction, as long-context question-answering data is relatively scarce on the internet. To address this, prevalent long-context benchmarks have utilized synthetic tasks like passkey retrieval [mohtashami2023landmark], needle-in-a-haystack (NIAH) [needleinhaystack, zhao2024longagent], and variable tracking [hsieh2024ruler] to evaluate long-context LLMs. However, these tasks are often unrealistic and involve reasoning processes that differ significantly from those in real-world applications. Alternatively, some research efforts have involved human annotation of realistic questions and gold answers over one or multiple long documents [bamboo, wang2024leave, li2023loogle]. However, creating realistic long-context tasks from extensive texts is both challenging and time-consuming, even for human experts [wang2024leave]. This limitation restricts the expansion of datasets to accommodate arbitrary context lengths and the ability to support controllable context. As shown in Table 1, existing benchmarks either rely on a limited number of synthetic tasks, demand significant human effort to read long contexts, or lacking controllable contexts and support for arbitrary context lengths. Furthermore, existing datasets [bamboo, wang2024leave, li2023loogle] often utilize documents from specific domains, such as financial reports or legal cases, as input, which can inherently limit the diversity of task categories. Consequently, they tend to focus on a narrow set of tasks, such as comparison or classification, rather than evaluating more complex and challenging tasks that require chain-of-thought reasoning.
@@ 2 Related Work
Long-Context Large Language Models Recent advancements in deep learning system have significantly propelled the development of long-context large language models (LLMs). One of the key challenges in scaling these models is the quadratic time and space complexity inherent in computing self-attention over long sequences. To mitigate this computational burden, efficient self-attention algorithms [dao2022flashattention, dao2023flashattention, liu2023ring] have been introduced, reducing memory overhead, and novel training methods [li2021sequence, chen2023longlora] facilitate the training of these long-context models. As Rotary Position Embedding (RoPE) [su2024roformer] is widely used for positional encoding in many open-source models [llama3-1, qwen2.5, mistrallarge2], recent research [pi, xiong2023effective, peng2023yarn, liu20242, ding2024longrope, pose] has focused on adapting RoPE from pre-trained short-context models to effectively handle longer sequences. Moreover, new architectures [mamba, rwkv, bulatov2022recurrent, rmt, botev2024recurrentgemma] have been developed to efficiently process long-context inputs. Consequently, state-of-the-art language models [openai2024gpt4o, openai2024gpt4omini, reid2024gemini, anthropic2024claud35sonnet, llama3-1, mistrallarge2, qwen2.5, glm4] now support context windows ranging from 128K to millions of tokens, enabling the exploration of reasoning abilities over extensive contexts with LLMs.
@@
Long-Context Benchmarks As the context window of current LLMs expands rapidly, numerous benchmarks have been proposed to evaluate their capabilities. In early benchmarks such as ZeroSCROLLS [zeroscrolls], L-Eval [leval], BAMBOO [bamboo], LongBench [bai2023longbench], and LooGLE [li2023loogle], the average input length remains under 25K tokens, which is far shorter than the context window size supported by existing LLMs. Recently, some research has begun to explore using synthetic datasets, which can support controllable context lengths, to evaluate the long-context abilities of LLMs. Needle-in-a-Haystack and its variants [needleinhaystack, zhao2024longagent] primarily evaluate retrieval abilities by inserting relevant information into extensive irrelevant corpora and testing the LLMs’ capacity to extract it. Additionally, RULER [hsieh2024ruler] constructs synthetic tasks based on code-like flexible configurations to assess LLM performance over long contexts. While synthetic tasks can support the evaluation of arbitrarily long contexts, they are limited in scope, focusing on a narrow set of tasks and failing to comprehensively evaluate the reasoning abilities of LLMs in realistic scenarios. Other benchmarks like InfiniteBench [zhang2024infty] and Loong [wang2024leave] use human annotations to create questions from given long texts, which contain more diverse tasks but are both time-consuming and costly. Our proposed benchmark, LongReason, focuses on evaluating the long-context reasoning abilities of LLMs, which are created automatically from short reasoning questions without heavy human effort in reading the long context. We conduct a detailed comparison with existing benchmarks in Table 1.
@@ 3.1 Long-context Reasoning Question Construction via Context Expansion
Long-Context Reasoning Question Construction Through Context Expansion Finally, we construct the long-context version of each question by embedding each passage $C_{e}^{i}$ from the expanded context $\bar{C}{\text{expanded}}$ at random positions within a set of irrelevant passages $\bar{C}{\text{irrelevant}}$, forming the final long-context reasoning questions. To create $\bar{C}{\text{irrelevant}}$ , we first collect passages from the Pile [gao2020pile] and use an LLM to rewrite each passage to minimize stylistic differences between the synthesized background passages and the irrelevant passages. These rewritten passages are then compiled to form the set of irrelevant passages $\bar{C}_{\text{irrelevant}}$. In LongReason, GPT-4 is used for all data synthesis and self-verification. For each question, we evaluate multiple versions of the synthesized question for comparison, including the original question $Q_{\text{short}}$, the expanded version $Q_{\text{expanded}}$, and long-context versions with context lengths ranging from $8K$ to $128K$. Furthermore, similar to NIAH [needleinhaystack], our pipeline is capable of generating reasoning questions with even longer contexts by incorporating additional irrelevant information.
@@ 3.2 The Statistics of LongReason
00footnotetext: Mistral Large 2 generates an empty response when the question context length reaches 128K.
@@ 4.1 Experimental setup
Models & Inference Setup We select a set of representative LLMs that support long context windows, including 6 closed-source models from 3 model families (GPT, Gemini and Claude) and 15 open-source models spanning a wide range of model sizes (3B to 123B) and claimed context lengths (8K to 2M). Detailed information about these models can be found in Appendix 6. For open-source models, we utilize vLLM [vllm], which enables efficient KV cache memory management during inference time. All inferences are performed using bfloat16 precision on 8 NVIDIA A100 GPUs with greedy decoding (temperature=0).
Evaluation setup We evaluate all models on LongReason, which comprises 794 questions, each featuring multiple variations, including the original version, expanded versions, and long-context versions with context lengths of 8K, 16K, 32K, 64K, and 128K. Each input is constructed using a predefined zero-shot chain-of-thought template that combines the background context, followed by the corresponding final inquiry. Detailed information about the prompt template is provided in Appendix 9. To assess the reasoning performance of the LLMs, we extract the predicted choice by identifying the first character following the phrase “the answer is” and compare it to the ground-truth option for accuracy.
@@ 6 Model Information
Model Aligned Size Context Length Huggingface [huggingface] / API
GPT-4o [openai2024gpt4o] - 128K gpt-4o-2024-08-06
GPT-4o-mini [openai2024gpt4omini] - 128K gpt-4o-mini-2024-07-18
Gemini-1.5-Pro [gemini] - 2M gemini-1.5-pro-002
Gemini-1.5-Flash [gemini] - 2M gemini-1.5-flash-002
Claude-3.5-Sonnet[anthropic2024claud35sonnet] - 200K claude-3-5-sonnet-20240620
Claude-3.5-Haiku[anthropic2024claud35sonnet] - 200K claude-3-5-haiku-20241022
Mistral-Large2 [mistrallarge2] 123B 128K mistralai/Mistral-Large-Instruct-2407
[…]
Mistral-Small [mistraltec] 22B 32K mistralai/Mistral-Small-Instruct-2409
Mistral-Nemo [mistralnemo] 12B 1M mistralai/Mistral-Nemo-Instruct-2407
Mistral-7B [mistral] 7B 32K mistralai/Mistral-7B-Instruct-v0.3
Llama3.1 [llama3-1] 70B 128K meta-llama/Meta-Llama-3.1-70B-Instruct
Llama3.1 [llama3-1] 8B 128K meta-llama/Meta-Llama-3.1-8B-Instruct
Qwen2.5 [qwen2.5] 72B 128K Qwen/Qwen2.5-72B-Instruct
Qwen2.5 [qwen2.5] 32B 128K Qwen/Qwen2.5-32B-Instruct
Qwen2.5 [qwen2.5] 14B 128K Qwen/Qwen2.5-14B-Instruct
Qwen2.5 [qwen2.5] 7B 128K Qwen/Qwen2.5-7B-Instruct
Qwen2.5 [qwen2.5] 3B 32K Qwen/Qwen2.5-3B-Instruct
GLM4-9B [glm4] 9B 128K THUDM/glm-4-9b-chat
[…]
Phi3.5-mini [phi3] 14B 128K microsoft/Phi-3.5-mini-instruct
@@
7 Hyperparameters for LongReason Construction
@@ 9 Prompts
Please expand the above material into multiple independent paragraphs with around 200 words in English, while meeting
the following requirements. 1. Ensure that every key piece of information from the material appears in one paragraph of
the expanded text. Try to place the key information in the middle of the expanded paragraphs.
2. The expanded material need to avoid introducing additional knowledge, reasoning, or any content that might influence
the answer to the "Question about the material."
3. The expanded text should clearly relate to the "Question about the material." Please include hints or references to the
question within each paragraph to maintain this connection. However, do not use words like "question" or "query"
explicitly in the expanded text.
4. Each paragraph should be a standalone piece of text, comprehensible without needing to refer to other paragraphs.
Minimize the use of pronouns, particularly those referring to other paragraphs.
@@
Please analyze how to appropriately add background information to expand the material into multiple independent
paragraphs based on the given requirements. Additionally, assess whether the provided material can be easily divided into
multiple paragraphs. If not please only provide only one paragraph in the expanded material.
@@
Present the expanded material as a series of independent paragraphs that meet the above requirements. Add the index
like "1." (do not use any format here) at the beginning of each paragraph (starting from 1), use English, and do not
include any extra information.
@@ 2. Does the expanded material contain information that will affect the answer to the question?
Please analyze the expanded material and compare it with the original material. Then, combine the question and analyze the
three questions above.
### Does the expanded material contain all the key information from the original material based on
the analysis?
@@
### Does the expanded material contain information that will affect the answer to the question based
on the analysis?
@@
### Are all the paragraphs in the expanded material are related to the main topic/character of the
question based on the analysis?